how much sand do you need to add to 10ft to the third powerof soil with 22% sand in order to make a soil with 35% sand?
1
Expert's answer
2012-09-27T10:31:04-0400
10 ft³ of soil with 22% sand contains

S = 10[ft³]*0.22 = 2.2 ft³

of sand. In order to make a soil with 35% sand we need to add M ft³ of sand that must satisfy following condition:

(10+M)*0.35 = M

3.5 + 0.35M = M

3.5 = 0.65M

M = 3.5/0.65 ft³ = 70/13 ft³.

So, we need to add 70/13 ft³ of sand.



&